If the ionisation potential of an atom .is 20V its 1st excitation potential will be

Grade:11

3 Answers

Vikas TU
7 years ago
Ionisation poetential is the poetential calculate from the orbit n=1 to the n = infinity and that is given as:
E = > 13.6*z^2
z^2 = 20/13.6............(1)
Excitaion energy is given as : => -13.6*(z^2/4 – z^2/1)
from eqn. (1) we get,
=> -13.6*(5/13.6 – 20/13.6)
=> 15 V
ved
6 years ago
Let us consider ionisation potential for hydrogen atomSo I.P.=13.6evNow converting 13.6ev into potential difference (v) we divide it by x.i.e. 13.6/x=20vTherfore x=13.6/20Now its first excitation potentail will be 13.6 * 3/4 ev Now converting it into potential diffrence (v)We divide it by x i.e.13.6/20=(13.6*3)/(4*x)=(13.6*3*20)/(13.6*4)=15vHence the answer....
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Ionisation poetential is the poetential calculate from the orbit n=1 to the n = infinity and that is given as:
E = > 13.6*z2
z2 = 20/13.6............(1)
Excitation energy is given as : => – 13.6*(z2/4 – z2/1)
from eqn. (1) we get,
=> – 13.6*(5/13.6 – 20/13.6)
=> 15 V
 
Thanks and regards,
Kushagra